OpenStudy (anonymous):

determine P(cos 225 degrees, sin 225 degrees). Then use those vaules to verify that x^2 = y^2 = 1

OpenStudy (anonymous):

The cos of 225 and sin of 225 are both \[\frac{ \sqrt{2} }{ 2 }\], so if x = y, then x^2 must = y^2.

OpenStudy (anonymous):

That square root is negative too.

OpenStudy (anonymous):

x=cos225=cos(180+45)=-cos45=-1/sqrt2 y=sin225=sin(180+45)=-sin45=-1/sqrt2 x^2=1/2 y^2=1/2 x^1/2=y^1/2